OpenStudy (eric_d):

The functions f and g are defined for x is element of real number (excluding -1,0,1) by

OpenStudy (eric_d):

f(x)= 1+x/1-x and g(x)=1/x

OpenStudy (eric_d):

@mathmale

OpenStudy (eric_d):